calculate the radioactive constant for an element whose half life is 20 years
1
Expert's answer
2020-09-28T08:06:03-0400

The radioactive constant

"\lambda=\frac{\ln 2}{t_{1/2}}\\
=\frac{\ln 2}{20\times 365\times 24\times 3600\:\rm s}=1.098\times 10^{-9}\:\rm s^{-1}"